# A new framework for nonlinear Kalman filters

### v6

@@ 5.2 Terrain-referenced navigation
5.3 Synchronous generator state estimation
@@
All the model parameters are selected to be the same as in [23]. The system’s state transition and measurement functions can be written as (27, 28). Note that the noises are omitted from the equations.
@@
The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$. The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{–4}$, $10^{–10}$, $10^{–4}$, and $10^{–4}$. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are $10^{–10}$, $10^{–16}$, $10^{–10}$, and $10^{–10}$.
5.4 Pendulum state estimation
As shown in Fig. 11(c), one end of the pendulum is fixed to the ceiling with a thin rope. The system’s states are the pendulum’s angular position $\theta$ and angular speed $\omega$. The measurement $y_{k}$ is the horizontal component of tension in the rope. The system input is the external torque exerted on the pendulum, which is zero in this example. The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is $g=9.8$m/s2. The time interval is set to $\Delta t=0.01$s. According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (29,30). Note that the noises are omitted in the equations.
@@ 5.4 Pendulum state estimation
5.5 Battery state estimation
This example is adapted from a research paper [24]. As shown in Fig. 11(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$. A battery’s two most important states are its state of charge (SOC) and state of health (SOH). The SOC is the ratio of the battery’s remaining capacity to its present maximum capacity, and the SOH is the ratio of the battery’s present maximum capacity to its initial maximum capacity. In general, SOC describes the battery’s energy level, while SOH describes the battery’s aging level. The states of this system are the SOC, SOH, and capacitor voltage $U_{c}$. The measurement $y_{k}$ is the battery’s terminal voltage. The input is the charging current, which, in this example, is a 3-level square wave shown below:
@@ 5.5 Battery state estimation
where $a_{100}=$ [1390.38, -6961.31, 14760.31, -17230.92, 12055.71, -5162.75, 1330.60, -196.37, 15.60, 2.96], $a_{80}=$ [813.94, -4229.96, 9345.49, -11415.38, 8396.15, -3801.07, 1043.09, -165.29, 14.28, 2.96]. The time interval is set to $\Delta t=1$ s, and the total simulation time is 180 seconds. The system’s state transition and measurement functions can be written as (33,34). Note that the noises are omitted in the equations.
@@
6 The detailed algorithm of various nonlinear Kalman filters used in the paper
The detailed algorithms of EKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 36. Note that all these algorithms degrade to the conventional framework when the ”recalibrate” and ”update” steps are replaced by (9).
@@ 6 The detailed algorithm of various nonlinear Kalman filters used in the paper
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
@@
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
@@
7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
8:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})$
9:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}+\boldsymbol{Q}_{k}$
@@
11:     $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k-1}}$
12:     $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{\hat{x}}_{k|k-1})$
13:     $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
14:     $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
15:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
16:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
17:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
@@
19:     $\boldsymbol{H}_{k|k}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k}}$
20:     $\boldsymbol{P}_{xy,k|k}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
21:     $\boldsymbol{P}_{y,k|k}=\boldsymbol{H}_{k|k}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
22:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
23:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
@@
25:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
26:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
27:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
@@
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, measurements number $n_{m}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
@@
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
@@
7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
8:     $\boldsymbol{F}_{i,k}^{*}=\big{[}\frac{\partial^{2}f_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}\big{]}_{mn},i=1,\ldots,n_{x}$
9:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{F}_{i,k}^{*}\boldsymbol{P}_{k-1|k-1})\big{]}_{i}$
10:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}$
11:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{F}_{i,k}^{*}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{j,k}^{*}\boldsymbol{P}_{k-1|k-1})\big{]}_{ij}+\boldsymbol{Q}_{k}$
@@
13:     $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k-1}}$
14:     $\boldsymbol{H}_{i,k|k-1}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k|k-1}}\big{]}_{mn},i=1,\ldots,n_{m}$
15:     $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k-1}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{i}$
16:     $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
17:     $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
18:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k-1}^{*}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{j,k|k-1}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{ij}$
19:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
20:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
21:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
@@
23:     $\boldsymbol{H}_{k|k}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k}}$
24:     $\boldsymbol{H}_{i,k|k}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k|k}}\big{]}_{mn},i=1,\ldots,n_{m}$
25:     $\boldsymbol{P}_{xy,k|k}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
26:     $\boldsymbol{P}_{y,k|k}=\boldsymbol{H}_{k|k}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
27:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k}^{*}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{j,k|k}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{ij}$
28:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
29:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
@@
31:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
32:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
33:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
@@
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$, hyperparameters $\alpha=10^{-3},\beta=2,\kappa=0$

### v10

@@ B.2 Terrain-referenced navigation
B.3 Synchronous generator state estimation
@@
All the model parameters are selected to be the same as in [27]. The system’s state transition and measurement functions can be written as (31, 32). Note that the noises are omitted from the equations.
@@
The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$. The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{-4}$, $10^{-10}$, $10^{-4}$, and $10^{-4}$. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are $10^{-10}$, $10^{-16}$, $10^{-10}$, and $10^{-10}$.
B.4 Pendulum state estimation
As shown in Fig. 11(c), one end of the pendulum is fixed to the ceiling with a thin rope. The system’s states are the pendulum’s angular position $\theta$ and angular speed $\omega$. The measurement $y_{k}$ is the horizontal component of tension in the rope. The system input is the external torque exerted on the pendulum, which is zero in this example. The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is $g=9.8$m/s2. The time interval is set to $\Delta t=0.01$s. According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (33,34). Note that the noises are omitted in the equations.
@@ B.4 Pendulum state estimation
B.5 Battery state estimation
This example is adapted from a research paper [28]. As shown in Fig. 11(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$. A battery’s two most important states are its state of charge (SOC) and state of health (SOH). The SOC is the ratio of the battery’s remaining capacity to its present maximum capacity, and the SOH is the ratio of the battery’s present maximum capacity to its initial maximum capacity. In general, SOC describes the battery’s energy level, while SOH describes the battery’s aging level. The states of this system are the SOC, SOH, and capacitor voltage $U_{c}$. The measurement $y_{k}$ is the battery’s terminal voltage. The input is the charging current, which, in this example, is a three-level square wave shown below:
@@ B.5 Battery state estimation
where $a_{100}=[1390.38,-6961.31,14760.31,-17230.92,$ 12055.71,5162.75,1330.60,196.37,15.60,2.96]12055.71,-5162.75,1330.60,-196.37,15.60,2.96]12055.71 , - 5162.75 , 1330.60 , - 196.37 , 15.60 , 2.96 ], $a_{80}=$ [813.94,4229.96,9345.49,11415.38,8396.15,3801.07,[813.94,-4229.96,9345.49,-11415.38,8396.15,-3801.07,[ 813.94 , - 4229.96 , 9345.49 , - 11415.38 , 8396.15 , - 3801.07 , 1043.09,165.29,14.28,2.96]1043.09,-165.29,14.28,2.96]1043.09 , - 165.29 , 14.28 , 2.96 ]. The time interval is set to $\Delta t=1$ s, and the total simulation time is 180 seconds. The system’s state transition and measurement functions can be written as (37,38). Note that the noises are omitted in the equations.
@@
Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
The detailed algorithms of EKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 35. Note that all these algorithms degrade to the conventional framework when the “Recalibrate” and “Back out” steps are replaced by (10).
@@ Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
@@
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
@@
7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
8:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
9:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
@@
11:     $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k-1}}$
12:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})$
13:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
14:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
15:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
16:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
17:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
@@
19:     $\bm{H}_{k|k}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k}}$
20:     $\bm{P}_{xy,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
21:     $\bm{P}_{y,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
22:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
23:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
@@
25:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
26:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
27:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
@@
1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, states number $n_{x}$, measurements number $n_{m}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
@@
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
@@
7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
8:     $\bm{F}_{i,k}^{*}=\big{[}\frac{\partial^{2}f_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}\big{]}_{mn},i=1,\ldots,n_{x}$
9:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})+\frac{1}{2}\big{[}\text{tr}(\bm{F}_{i,k}^{*}\bm{P}_{k-1|k-1})\big{]}_{i}$
10:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}$
11:      $+\frac{1}{2}\big{[}\text{tr}(\bm{F}_{i,k}^{*}\bm{P}_{k-1|k-1}\bm{F}_{j,k}^{*}\bm{P}_{k-1|k-1})\big{]}_{ij}+\bm{Q}_{k}$
@@
13:     $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k-1}}$
14:     $\bm{H}_{i,k|k-1}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k|k-1}}\big{]}_{mn},i=1,\ldots,n_{m}$
15:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{i}$
16:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
17:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
18:      $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
19:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
20:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
21:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
@@
23:     $\bm{H}_{k|k}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k}}$
24:     $\bm{H}_{i,k|k}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k|k}}\big{]}_{mn},i=1,\ldots,n_{m}$
25:     $\bm{P}_{xy,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
26:     $\bm{P}_{y,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
27:       $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
28:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
29:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
@@
31:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
32:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
33:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
@@
Algorithm 5 The cubature Kalman filter with the new framework
1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, states number $n_{x}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
2:Initialization:
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$,
8:        where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
9:     $\bm{\mathcal{X}}_{i,k-1}=\bm{\hat{x}}_{k-1|k-1}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
10:     $\bm{\mathcal{X}}_{i+n_{x},k-1}=\bm{\hat{x}}_{k-1|k-1}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
11:     $\bm{\mathcal{X}}_{i,k|k-1}^{*}=\bm{f}(\bm{\mathcal{X}}_{i,k-1},\bm{u}_{k-1}),i=1,\ldots,2n_{x}$
12:     $\bm{\hat{x}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{X}}_{i,k|k-1}^{*}$
13:     $\bm{P}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}^{*}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{X}}_{i,k|k-1}^{*}$
14:        $-\bm{\hat{x}}_{k|k-1}]^{T}+\bm{Q}_{k}$
15:    Update:
16:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k|k-1})$
17:     $\bm{\mathcal{X}}_{i,k|k-1}=\bm{\hat{x}}_{k|k-1}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
18:     $\bm{\mathcal{X}}_{i+n_{x},k|k-1}=\bm{\hat{x}}_{k|k-1}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
19:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
20:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k-1}$
21:     $\bm{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
22:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
23:     $\bm{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
24:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
25:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
26:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
27:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
28:    Recalibrate:
29:     $\bm{\mathcal{X}}_{i,k|k}=\bm{\hat{x}}_{k|k}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
30:     $\bm{\mathcal{X}}_{i+n_{x},k|k}=\bm{\hat{x}}_{k|k}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
31:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
32:     $\bm{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k}$
33:     $\bm{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
34:     $\bm{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
35:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
36:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
37:    Back out:
38:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
39:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
40:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
41:     end if
42:end for
@@
1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, states number $n_{x}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$, hyperparameters $\alpha=10^{-3},\beta=2,\kappa=0$